# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v1

@@ C.1 Self-energies
$\begin{split}&\Sigma_{\mathrm{BF}}=\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\frac{d\Omega}{2\pi}\left[\right.T\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{0}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)\\ &\hskip 88.2037pt-\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{00}(\textbf{P}-\textbf{k},\Omega-\omega)\left.\right]\\ &+\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\frac{d\Omega}{2\pi}\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{00}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)e^{i\Omega 0^{+}},\end{split}$ (74)
where $G_{\mathrm{F}}^{00}$ is a bare fermionic Green’s function with $\mu_{\mathrm{F}}^{0}=0^{-}$. The second frequency integral is evaluated via contour integration closing on the left-hand side and contributing with a pole and a branch cut when they fall in this region. Instead, the first frequency integration is split in three regions: $[-\infty,-\Omega_{c}]\cup[-\Omega_{c},\Omega_{c}]\cup[\Omega_{c},\infty]$. For $|\Omega|\geq\Omega_{c}$ the integral is carried out analytically exploiting the subleading asymptotic expression of the integrand. The remaining frequency range is instead treated numerically.
Regarding integration in momentum space, the angular integral can be carried out analytically, since the dependence on the angle appears only in the bare Green’s function. Finally, radial integrals are evaluated numerically up to a natural cut-off given by the position of the highest Fermi step in the integrand function. In particular, we find one step arising from the pole of the $T$-matrix and up to two steps coming from the poles of the fermionic Green’s function.
@@ C.2 Momentum distributions and densities
${n_{\mathrm{F}}}\left(\textbf{k}\right)$ $=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}{G_{\mathrm{F}}}\left(\textbf{k},\omega\right)e^{i\omega 0^{+}}$
$\simeq 2\int_{0}^{\omega_{c}}\frac{d\omega}{2\pi}\text{Re}\left[{G_{\mathrm{F}}}\left(\textbf{k},\omega\right)-G_{\mathrm{F}}^{0}\left(\textbf{k},\omega\right)\right]+\Theta\left(-\xi^{\mathrm{F}}_{{\mathbf{k}}}\right),$ (75)
@@
$\simeq 2\;\int_{0}^{\omega_{c}}\frac{d\omega}{2\pi}\text{Re}\left[{G_{\mathrm{B}}}\left(\textbf{k},\omega\right)-{G_{\mathrm{B}}^{0}}^{\prime}\left(\textbf{k},\omega\right)\right]$
$\quad\quad+\frac{1}{2}\left[\frac{\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}}{\sqrt{\left(\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}\right)^{2}-\Sigma^{2}_{12}}}-1\right],$ (76)
@@
${G_{\mathrm{B}}^{0}}^{\prime}\left(\textbf{k},\omega\right)=\left(i\omega-\xi^{\mathrm{B}}_{\textbf{k}}-\Sigma_{11}+\frac{\Sigma_{12}^{2}}{i\omega+\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}}\right)^{-1}.$ (77)
@@
${\Sigma_{\mathrm{F}}}$ $\left(\textbf{k},\omega\right)\xrightarrow{\omega\to\infty}n_{0}\Gamma\left(\textbf{k},\omega\right)-\left({{C_{\mathrm{BF}}}\over 4m_{r}^{2}}\right)G_{\mathrm{B}}^{0}\left(\textbf{k},-\omega\right),$ (78)
${\Sigma_{\mathrm{B}}}$ $\left(\textbf{k},\omega\right)\xrightarrow{\omega\to\infty}\Sigma_{11}+n_{{\mu_{\mathrm{F}}}}T\left(\textbf{k},\omega\right)-\left({{C_{\mathrm{BF}}}\over 4m_{r}^{2}}\right)G_{\mathrm{F}}^{0}\left(\textbf{k},-\omega\right)$ (79)
The last terms in Eq. (75) and in Eq. (76) originate from the frequency integration of the added and subtracted Green’s functions, $G_{\mathrm{F}}^{0}({\mathbf{k}},\omega)$ and ${G_{\mathrm{B}}^{0}}^{\prime}({\mathbf{k}},\omega)$ respectively.
@@
${n_{\mathrm{F}}}$ $=\int_{0}^{\infty}\frac{dk}{2\pi}k\>{n_{\mathrm{F}}}\left(k\right)$
$=\int_{0}^{k_{c}}\frac{dk}{2\pi}k\>{n_{\mathrm{F}}}\left(k\right)+\frac{\frac{{C_{\mathrm{BF}}}}{4m_{r}}}{2\pi\left(\frac{k^{2}_{c}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)},$ (80)
${n_{\mathrm{B}}}$ $=n_{0}+\int_{0}^{\infty}\frac{dk}{2\pi}k\>{n_{\mathrm{B}}}\left(k\right)=n_{0}+\int_{0}^{k_{c}}\frac{dk}{2\pi}k\>{n_{\mathrm{B}}}\left(k\right)$
$\quad\quad\quad+\frac{\frac{{C_{\mathrm{BF}}}}{4m_{r}}}{2\pi\left(\frac{k^{2}_{c}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)}+\frac{m_{\mathrm{B}}\Sigma^{2}_{12}}{8\pi\left(\frac{k_{c}^{2}}{2m_{\mathrm{B}}}-{\mu_{\mathrm{B}}}\right)},$ (81)
@@ C.3 BF contact parameter ${C_{\mathrm{BF}}}$
${{C_{\mathrm{BF}}}\over 4m_{r}^{2}}$ $=\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\frac{d\Omega}{2\pi}T\left(\textbf{P},\Omega\right)e^{i\Omega 0^{+}}$ (82)
$=\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\frac{d\Omega}{2\pi}\left[T\left(\textbf{P},\Omega\right)-\Gamma^{\mathrm{SC}}\left(\textbf{P},\Omega\right)\right]+$
$\hskip 42.67912pt\int\frac{d{\mathbf{P}}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\frac{d\Omega}{2\pi}\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)e^{i\Omega 0^{+}}$ (83)
$=\int\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\left\{2\text{Re}\left[\int_{0}^{\Omega_{c}}\frac{d\Omega}{2\pi}\left[T\left(\textbf{P},\Omega\right)-\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)\right]\right.\right.+$
$\quad\quad\quad\left.\frac{2\pi\left(n_{0}+n_{{\mu_{\mathrm{F}}}}\right)}{m_{r}^{2}}\frac{\pi/2-i\left(\ln\Omega_{c}-\ln{\varepsilon_{0}}\right)}{\pi^{2}/4+\ln^{2}\Omega_{c}-\ln^{2}{\varepsilon_{0}}}\right]+$
$\frac{2\pi}{m_{r}}$ $\left.\left[{\varepsilon_{0}}\Theta\left(-\xi^{\mathrm{\mathrm{CF}}}_{\textbf{P}}\right)+\Theta(-z_{c})\int_{z_{c}}^{0}\frac{dx}{\ln^{2}\left(x-z_{c}\right)-\ln^{2}{\varepsilon_{0}}+\pi^{2}}\right]\right\}$ (84)
where $\Omega_{c}$ is the high-frequency cutoff, the term in the second line of Eq. (84) comes from the analytical integration of the asymptotic form of $T({\mathbf{P}},\Omega)-\Gamma_{\mathrm{SC}}({\mathbf{P}},\Omega)$ in the range $\left[\Omega_{c},+\infty\right]$ with the definition $n_{{\mu_{\mathrm{F}}}}=2{m_{\mathrm{F}}}{\mu_{\mathrm{F}}}\Theta({\mu_{\mathrm{F}}})$, whereas in the last line $z_{c}=P^{2}/2M-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}$ and $\xi_{\textbf{P}}^{\mathrm{CF}}=P^{2}/2M-{\mu_{\mathrm{CF}}}$. The remaining integrations over $\Omega$, $x$ and $P$ are evaluated numerically.
Appendix D Analytic evaluations in the strong-coupling limit
The integral in expression (III.1) can be computed analytically and is given in closed form as
$\begin{split}\frac{n_{\mathrm{CF}}}{n_{\mathrm{F}}}=&\,\Theta(y_{+})\left[{y_{+}\over 2}-\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y_{+}\over 2}\right)^{2}}\right.\\ &\phantom{aaaaaaaaaa}+\left.\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}\right)^{2}}\right]\\ +&\,\Theta(y_{-})\left[{y_{-}\over 2}+\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y_{-}\over 2}\right)^{2}}\right.\\ &\phantom{aaaaaaaaaa}-\left.\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}\right)^{2}}\right],\end{split}$ (85)
@@ Appendix D Analytic evaluations in the strong-coupling limit
$\bar{\tilde{\xi}}^{\mathrm{CF}}_{y}+\bar{\xi}^{\mathrm{F}}_{y}\pm\sqrt{(\bar{}\xi^{\mathrm{CF}}_{y}-\bar{\xi}^{\mathrm{F}}_{y})^{2}+4\bar{\Delta}_{0}^{2}}=0,$ (86)
where $\bar{\tilde{\xi}}_{y}^{\mathrm{CF}}=m_{\rm F}y/M-\bar{\tilde{\mu}}_{\mathrm{CF}}$ and $\tilde{\xi}_{y}^{\mathrm{F}}=y-\bar{\mu}_{\rm F}$, and all quantities with a bar on top are in units of $E_{\rm F}$.
The Hugenholtz-Pines condition (III.4) can be evaluated analytically and, in terms of the quantity
$I_{\pm}(y)=\ln\left[\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y\over 2}\right)^{2}}\pm\left({\tilde{\mu}_{\mathrm{CF}}}-\bar{\mu}_{\rm F}+{y\over 2}\right)\right],$ (87)
@@
${\mu_{\mathrm{B}}}=-4{\varepsilon_{0}}\Big{\{}[I_{-}(y_{+})-I_{-}(0)]\Theta(y_{+})+[I_{+}(y_{-})-I_{+}(0)]\Theta(y_{-})\Big{\}}$ (88)
with ${y_{+}}$ and ${y_{-}}$ defined as in Eq. 85.
In the special case $x=1$ and for $g\to\infty$, $\Delta_{0}\to 0$ (as shown in Fig. 14) and the above expression acquires the simplified and recursive form
${\mu_{\mathrm{B}}}=-4{\varepsilon_{0}}\ln\left[1+\frac{{\mu_{\mathrm{CF}}}}{{\mu_{\mathrm{B}}}+{\varepsilon_{0}}}\right],$ (89)
@@
where ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}$ since $\Sigma_{\mathrm{CF}}=0$ and the definition ${\mu_{\mathrm{CF}}}={\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ has been used.
@@
With the replacement $t\equiv({\mu_{\mathrm{B}}}+{\varepsilon_{0}})/{\mu_{\mathrm{CF}}}$ it holds that ${\mu_{\mathrm{B}}}/{\varepsilon_{0}}=t{\mu_{\mathrm{CF}}}/{\varepsilon_{0}}-1$. To leading order in the small parameter ${\mu_{\mathrm{CF}}}/{\varepsilon_{0}}$, Eq. 89 reduces to ${1\over 4}=\ln\left[1+\frac{1}{t}\right]$, with solution
$t\equiv{{\mu_{\mathrm{B}}}+{\varepsilon_{0}}\over{\mu_{\mathrm{CF}}}}=\frac{1}{\mathrm{e}^{1\over 4}-1}.$ (90)
@@
For balanced populations all bosons are expected to bind with fermions into pairs which implies that $n_{0}=0$ and, since $n_{\mathrm{CF}}=n_{\mathrm{F}}$, it must hold that ${\mu_{\mathrm{CF}}}=E_{\mathrm{F}}/2$ (as one can verify using expression (III.1)) allowing us to analytically determine the values of both the fermionic and bosonic chemical potentials through Eq. (90):

### v2

@@ C.1 Self-energies
$\Sigma_{\mathrm{BF}}$ $=\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\!\frac{d\Omega}{2\pi}\left[T\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{0}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)-\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{00}(\textbf{P}-\textbf{k},\Omega-\omega)\right]$
$+\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\!\frac{d\Omega}{2\pi}\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)G_{\mathrm{F}}^{00}\left(\textbf{P}-\textbf{k},\Omega-\omega\right)e^{i\Omega 0^{+}},$ (C.1)
where $G_{\mathrm{F}}^{00}$ is a bare fermionic Green’s function with $\mu_{\mathrm{F}}^{0}=0^{-}$. The second frequency integral is evaluated via contour integration closing on the left-hand side and contributing with a pole and a branch cut when they fall in this region. Instead, the first frequency integration is split into three regions: $[-\infty,-\Omega_{c}]\cup[-\Omega_{c},\Omega_{c}]\cup[\Omega_{c},\infty]$. For $|\Omega|\geq\Omega_{c}$ the integral is carried out analytically exploiting the subleading asymptotic expression of the integrand. The remaining frequency range is instead treated numerically.
Regarding integration in momentum space, the angular integral can be carried out analytically, since the dependence on the angle appears only in the bare Green’s function. Finally, radial integrals are evaluated numerically up to a natural cut-off given by the position of the highest Fermi step in the integrand function. In particular, we find one step arising from the pole of the many-body $T$-matrix and up to two steps coming from the poles of the fermionic Green’s function.
@@ C.2 Momentum distributions and densities
${n_{\mathrm{F}}}\left(\textbf{k}\right)$ $=\int_{-\infty}^{\infty}\!\frac{d\omega}{2\pi}{G_{\mathrm{F}}}\left(\textbf{k},\omega\right)e^{i\omega 0^{+}}\simeq 2\int_{0}^{\omega_{c}}\!\frac{d\omega}{2\pi}\text{Re}\left[{G_{\mathrm{F}}}\left(\textbf{k},\omega\right)-G_{\mathrm{F}}^{0}\left(\textbf{k},\omega\right)\right]+\Theta\left(-\xi^{\mathrm{F}}_{{\mathbf{k}}}\right),$ (C.2)
@@
$\simeq 2\int_{0}^{\omega_{c}}\!\frac{d\omega}{2\pi}\text{Re}\left[{G_{\mathrm{B}}}\left(\textbf{k},\omega\right)-{G_{\mathrm{B}}^{0}}^{\prime}\left(\textbf{k},\omega\right)\right]+\frac{1}{2}\left[\frac{\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}}{\sqrt{\left(\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}\right)^{2}-\Sigma^{2}_{12}}}-1\right],$ (C.3)
@@
${G_{\mathrm{B}}^{0}}^{\prime}\left(\textbf{k},\omega\right)=\left(i\omega-\xi^{\mathrm{B}}_{\textbf{k}}-\Sigma_{11}+\frac{\Sigma_{12}^{2}}{i\omega+\xi^{\mathrm{B}}_{\textbf{k}}+\Sigma_{11}}\right)^{-1}.$ (C.4)
@@
${\Sigma_{\mathrm{F}}}$ $\left(\textbf{k},\omega\right)\xrightarrow{\omega\to\infty}n_{0}\Gamma\left(\textbf{k},\omega\right)-\left({{C_{\mathrm{BF}}}\over 4m_{r}^{2}}\right)G_{\mathrm{B}}^{0}\left(\textbf{k},-\omega\right)$ (C.5)
${\Sigma_{\mathrm{B}}}$ $\left(\textbf{k},\omega\right)\xrightarrow{\omega\to\infty}\Sigma_{11}+n_{{\mu_{\mathrm{F}}}}T\left(\textbf{k},\omega\right)-\left({{C_{\mathrm{BF}}}\over 4m_{r}^{2}}\right)G_{\mathrm{F}}^{0}\left(\textbf{k},-\omega\right).$ (C.6)
The last terms in Eq. (C.2) and in Eq. (C.3) originate from the frequency integration of the added and subtracted Green’s functions, $G_{\mathrm{F}}^{0}({\mathbf{k}},\omega)$ and ${G_{\mathrm{B}}^{0}}^{\prime}({\mathbf{k}},\omega)$ respectively.
@@
${n_{\mathrm{F}}}$ $=\int_{0}^{\infty}\!\!\!\frac{dk}{2\pi}k\>{n_{\mathrm{F}}}\left(k\right)=\int_{0}^{k_{c}}\!\!\!\frac{dk}{2\pi}k\>{n_{\mathrm{F}}}\left(k\right)+\frac{\frac{{C_{\mathrm{BF}}}}{4m_{r}}}{2\pi\left(\frac{k^{2}_{c}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)},$ (C.7)
${n_{\mathrm{B}}}$ $=n_{0}+\int_{0}^{\infty}\!\!\!\frac{dk}{2\pi}k\>{n_{\mathrm{B}}}\left(k\right)$
$=n_{0}+\int_{0}^{k_{c}}\!\!\!\frac{dk}{2\pi}k\>{n_{\mathrm{B}}}\left(k\right)+\frac{\frac{{C_{\mathrm{BF}}}}{4m_{r}}}{2\pi\left(\frac{k^{2}_{c}}{2m_{r}}-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}\right)}+\frac{m_{\mathrm{B}}\Sigma^{2}_{12}}{8\pi\left(\frac{k_{c}^{2}}{2m_{\mathrm{B}}}-{\mu_{\mathrm{B}}}\right)},$ (C.8)
@@ C.3 BF contact parameter ${C_{\mathrm{BF}}}$
${{C_{\mathrm{BF}}}\over 4m_{r}^{2}}$ $=\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\!\frac{d\Omega}{2\pi}T\left(\textbf{P},\Omega\right)e^{i\Omega 0^{+}}$ (C.9)
$=\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\!\frac{d\Omega}{2\pi}\left[T\left(\textbf{P},\Omega\right)-\Gamma^{\mathrm{SC}}\left(\textbf{P},\Omega\right)\right]+\int\!\!\!\frac{d{\mathbf{P}}}{\left(2\pi\right)^{2}}\int_{-\infty}^{\infty}\!\frac{d\Omega}{2\pi}\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)e^{i\Omega 0^{+}}$ (C.10)
$=2\int\!\!\!\frac{d\textbf{P}}{\left(2\pi\right)^{2}}\left\{\text{Re}\left[\int_{0}^{\Omega_{c}}\!\frac{d\Omega}{2\pi}\big{[}T\left(\textbf{P},\Omega\right)-\Gamma_{\mathrm{SC}}\left(\textbf{P},\Omega\right)\big{]}\right.\right.$
$\hskip 71.13188pt+\left.\frac{2\pi\left(n_{0}+n_{{\mu_{\mathrm{F}}}}\right)}{m_{r}^{2}}\frac{\pi/2-i\left(\ln\Omega_{c}-\ln{\varepsilon_{0}}\right)}{\pi^{2}/4+\ln^{2}\Omega_{c}-\ln^{2}{\varepsilon_{0}}}\right]$
$\hskip 71.13188pt+\frac{2\pi}{m_{r}}\left.\left[{\varepsilon_{0}}\Theta\left(-\xi^{\mathrm{\mathrm{CF}}}_{\textbf{P}}\right)+\Theta(-z_{c})\int_{z_{c}}^{0}\frac{dx}{\ln^{2}\left(x-z_{c}\right)-\ln^{2}{\varepsilon_{0}}+\pi^{2}}\right]\right\}$ (C.11)
where $\Omega_{c}$ is the high-frequency cutoff, the term in the second line of Eq. (C.11) comes from the analytical integration of the asymptotic form of $T({\mathbf{P}},\Omega)-\Gamma_{\mathrm{SC}}({\mathbf{P}},\Omega)$ in the range $\left[\Omega_{c},+\infty\right]$ with the definition $n_{{\mu_{\mathrm{F}}}}=2{m_{\mathrm{F}}}{\mu_{\mathrm{F}}}\Theta({\mu_{\mathrm{F}}})$, whereas in the last line $z_{c}=P^{2}/2M-{\mu_{\mathrm{B}}}-{\mu_{\mathrm{F}}}$ and $\xi_{\textbf{P}}^{\mathrm{CF}}=P^{2}/2M-{\mu_{\mathrm{CF}}}$. The remaining integrals over $\Omega$, $x$ and $P$ are evaluated numerically.
Appendix D Analytic expressions in the strong-coupling limit
D.1 Composite-fermion density and Hugenholtz-Pines condition
The integral in expression (33) can be computed analytically, yielding
$\frac{n_{\mathrm{CF}}}{n_{\mathrm{F}}}=$ $\,\Theta(y_{+})\left[{y_{+}\over 2}-\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y_{+}\over 2}\right)^{2}}\right.+\left.\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}\right)^{2}}\right]$
$+$ $\,\Theta(y_{-})\left[{y_{-}\over 2}+\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y_{-}\over 2}\right)^{2}}\right.-\left.\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}\right)^{2}}\right],$ (D.1)
@@ D.1 Composite-fermion density and Hugenholtz-Pines condition
$\bar{\tilde{\xi}}^{\mathrm{CF}}_{y}+\bar{\xi}^{\mathrm{F}}_{y}\pm\sqrt{(\bar{\tilde{\xi}}^{\mathrm{CF}}_{y}-\bar{\xi}^{\mathrm{F}}_{y})^{2}+4\bar{\Delta}_{0}^{2}}=0.$ (D.2)
Here $\bar{\tilde{\xi}}_{y}^{\mathrm{CF}}=m_{\rm F}y/M-\bar{\tilde{\mu}}_{\mathrm{CF}}$ and $\tilde{\xi}_{y}^{\mathrm{F}}=y-\bar{\mu}_{\rm F}$, and all quantities with a bar on top are in units of $E_{\rm F}$.
The Hugenholtz-Pines condition (47) can be evaluated analytically and, in terms of the quantity
$I_{\pm}(y)=\ln\left[\sqrt{4\bar{\Delta}_{0}^{2}+\left(\bar{\tilde{\mu}}_{\mathrm{CF}}-\bar{\mu}_{\rm F}+{y\over 2}\right)^{2}}\pm\left({\tilde{\mu}_{\mathrm{CF}}}-\bar{\mu}_{\rm F}+{y\over 2}\right)\right],$ (D.3)
@@
${\mu_{\mathrm{B}}}=-4{\varepsilon_{0}}\Big{\{}\big{[}I_{-}(y_{+})-I_{-}(0)\big{]}\Theta(y_{+})+\big{[}I_{+}(y_{-})-I_{+}(0)\big{]}\Theta(y_{-})\Big{\}}.$ (D.4)
In the special case $x=1$ and for $g\to\infty$, one has $\Delta_{0}\to 0$ (as it can be seen in Fig. 14) and Eq. (D.4) can be simplified as
${\mu_{\mathrm{B}}}=-4{\varepsilon_{0}}\ln\left[1+\frac{{\mu_{\mathrm{CF}}}}{{\mu_{\mathrm{B}}}+{\varepsilon_{0}}}\right],$ (D.5)
@@
where ${\tilde{\mu}_{\mathrm{CF}}}={\mu_{\mathrm{CF}}}$ since $\Sigma^{0}_{\mathrm{CF}}=0$ for $x=1$, and the definition ${\mu_{\mathrm{CF}}}={\mu_{\mathrm{F}}}+{\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ has been used.
@@
By introducing $t\equiv({\mu_{\mathrm{B}}}+{\varepsilon_{0}})/{\mu_{\mathrm{CF}}}$, one has ${\mu_{\mathrm{B}}}/{\varepsilon_{0}}=t{\mu_{\mathrm{CF}}}/{\varepsilon_{0}}-1$. To leading order in the small parameter ${\mu_{\mathrm{CF}}}/{\varepsilon_{0}}$, Eq. (D.5) reduces to ${1\over 4}=\ln\left[1+\frac{1}{t}\right]$, with solution
$t\equiv{{\mu_{\mathrm{B}}}+{\varepsilon_{0}}\over{\mu_{\mathrm{CF}}}}=\frac{1}{\mathrm{e}^{1\over 4}-1}.$ (D.6)
@@
For balanced populations, all bosons are expected to bind with fermions into pairs. This implies $n_{0}=0$ and, since $n_{\mathrm{CF}}=n_{\mathrm{F}}$, ${\mu_{\mathrm{CF}}}=E_{\mathrm{F}}/2$ (as one can verify using expression (33)). Equation  (D.6), together with ${\mu_{\mathrm{CF}}}=E_{\mathrm{F}}/2$, analytically determine the values of both the fermionic and bosonic chemical potentials:
${\mu_{\mathrm{B}}}$ $=$ $-{\varepsilon_{0}}+\frac{1}{\mathrm{e}^{1\over 4}-1}{{E_{\mathrm{F}}}\over 2}$ (D.7)
${\mu_{\mathrm{F}}}$ $=$ $\frac{\mathrm{e}^{1\over 4}-2}{\mathrm{e}^{1\over 4}-1}{{E_{\mathrm{F}}}\over 2}.$ (D.8)
D.2 Fermion momentum distribution for small hybridization energy $\Delta_{0}$
We now derive an analytic expression for the fermion momentum distribution in the strong-coupling limit when the hybridization energy $\Delta_{0}\ll{E_{\mathrm{F}}}$. Within our theoretical approach, this condition occurs either when $x\to 1$ (with the condensate density depleted by the formation of molecules) or when $x\to 0$ (with the condensate density small because of the small boson number). For intermediate concentrations, $\Delta_{0}$ remains instead of order $E_{\rm F}$ in the strong-coupling limit (see Fig. 14). This analysis will provide us with a clear decomposition of $n_{\mathrm{F}}({\mathbf{k}})$ into its atomic and molecular contributions.
By expanding $G_{\mathrm{F}}({\mathbf{k}},\omega)$ to first order in $\Delta_{0}/{E_{\mathrm{F}}}$, one obtains
$G_{\mathrm{F}}({\mathbf{k}},\omega)\approx{1\over i\omega-\xi_{\mathbf{k}}^{\rm F}-{\Delta^{2}_{\mathrm{CF}}\over i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}}}+\frac{\Delta_{0}^{2}}{\left(i\omega-\tilde{\xi}_{\mathbf{k}}^{\mathrm{CF}}\right)\left(i\omega-\xi^{\mathrm{F}}_{\mathbf{k}}-{2\pi n_{\mathrm{CF}}\over m_{r}}\right)^{2}},$ (D.9)
where we have further approximated
$\frac{\Delta_{\mathrm{CF}}^{2}}{i\omega+\xi^{\mathrm{B}}_{{\mathbf{k}}}+\mu_{\mathrm{CF}}}\approx-\frac{\Delta_{\mathrm{CF}}^{2}}{\mu_{\mathrm{B}}}={2\pi n_{\mathrm{CF}}\over m_{r}}\equiv\Sigma_{\mathrm{F}}^{0}$ (D.10)
in the denominator of the second term. The quantity $2\pi n_{\mathrm{CF}}/m_{r}\equiv\Sigma_{\mathrm{F}}^{0}$ is analogous to $\Sigma^{0}_{\mathrm{CF}}$ of Sec. 3.1 in that it represents the repulsion induced by the mean-field of the molecules onto the unpaired fermions and can be interpreted as a Hartree mean-field shift of the chemical potential (${\tilde{\mu}_{\mathrm{F}}}\equiv{\mu_{\mathrm{F}}}-\Sigma_{\mathrm{F}}^{0}$).
Let us first consider the first term in Eq. (D.9). It can be recast in the more familiar (BCS-like) form
$\frac{\bar{u}^{2}_{k}}{i\omega-\bar{E}_{\mathbf{k}}^{+}}+\frac{\bar{v}^{2}_{k}}{i\omega-\bar{E}_{\mathbf{k}}^{-}}$ (D.11)
with
$\bar{E}_{\mathbf{k}}^{\pm}={1\over 2}\left(\xi_{\mathbf{k}}^{\mathrm{F}}-\xi^{\mathrm{B}}_{{\mathbf{k}}}\pm\sqrt{(\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}})^{2}+4\Delta_{\mathrm{CF}}^{2}}\right)$ (D.12)
and
$\bar{u}^{2}_{k}={1\over 2}\left(1+\frac{\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}}}{\sqrt{(\xi_{\mathbf{k}}^{\rm F}+\xi^{\mathrm{B}}_{{\mathbf{k}}})^{2}+4\Delta_{\mathrm{CF}}^{2}}}\right),\quad\quad\bar{v}^{2}_{k}=1-\bar{u}^{2}_{k}.$ (D.13)
$\bar{E}_{\mathbf{k}}^{+}\simeq\tilde{\xi}_{\mathbf{k}}^{\mathrm{F}},\quad\quad\quad\bar{E}_{\mathbf{k}}^{-}\simeq-\xi_{\mathbf{k}}^{\mathrm{B}},$ (D.14)